OpenStudy (anonymous):

Heather's family bought tickets at a local carnival. Her brother took 1/5 the number of tickets that her cousin took, which was half the number Heather took. If her brother took ten tickets, how many tickets did Heather take? 10 25 50 100

OpenStudy (amistre64):

Her brother took 1/5 the number of tickets that her cousin took b = c/5 which was half the number Heather took. c = h/2 If her brother took ten tickets, how many tickets did Heather take? if b=10, what is h ?

OpenStudy (amistre64):

b = c/5 ; c = 5b 5b = h/2; h = 10b

OpenStudy (anonymous):

@amistre64 ty!